# Tight contact structures without symplectic fillings are everywhere

### 27 Feb 2024 v4

@@ 2.2. ADC Bourgeois contact manifolds
In this section, we will assume that the first Chern class $c_{1}(\xi)$ of all contact structures are torsion. In this case there is a well-defined (rational) Conley-Zehnder index $\mu_{CZ}(\gamma)$ for any non-degenerate contractible periodic Reeb orbit $\gamma$. If the contact manifold is of dimension $12n-$ 1, then the SFT degree of $\gamma$ is defined as $|\gamma|\coloneqq n-3+\mu_{CZ}(\gamma)$. Notice that this does not correspond to the degree in symplectic cohomology of $\gamma$ when viewed as a periodic Hamiltonian orbit, which is just given by $n$ minus the Conley-Zehnder index.
We now recall the following definition from [Zho21b], that generalizes the one from [Laz20a]:
@@
A contact structure $(M,\xi)$ is called $k$ asymptotically dynamically convex ($k$-ADC) if there is a sequence of contact forms $\alpha_{1}\geq\alpha_{2}\geq\cdots\geq\alpha_{i}\geq\dots$ so that all contractible periodic orbits of $\alpha_{i}-$action $\leq D_{i}$ are non-degenerate and have degree $>k$, with $D_{i}\to\infty$.
In particular, 00-ADC as in [Zho21b] is just ADC as in [Laz20a]. An important special case is when there is a contact form $\alpha$ that is index-positive, meaning, following [CO18, Section 9.5], that all contractible periodic orbits are non-degenerate and have positive degree for $\alpha_{i}\equiv\alpha$ the given contact form. A very useful aspect of the ADC condition is that it is preserved under flexible surgery [Laz20a], whereas index-positivity is not.
@@
Consider the $12n-$ 1-dimensional Brieskorn spheres $\Sigma_{n}(k,2,\cdots,2)$ given as the link of the $A_{k-1}$-singularity, i.e.
@@
This is a contact manifold which is index-positive. The degrees of the generators are at least $42n-$ 4, provided that $n\geq 3$. This follows from [Ust99] in case $n=2m+1$ is odd, and from [vK08, Section 3] in general.
@@
for odd primes $p_{i}$, which according to [vK08, Section 4.1], satisfy the same properties in terms of their Conley-Zehnder indices. This also follows from a deep theorem of McLean [McL16, Theorem 1.1], which implies that the Brieskorn manifold $\Sigma_{n}(p_{1},p_{2},\cdots,p_{n+1})$ is ADC if the singularity is terminal.
@@
For the case of $3$-dimensional Brieskorn manifolds, the computations are more involved as the manifolds are not simply connected. One can view links of $A_{k-1}$-singularities as quotients of the standard contact $3$-sphere so that contractible periodic orbits on the links can be lifted to the $3$-sphere, which then have Conley-Zehnder indices at least $3$. However, when we construct a $7$-dimensional Bourgeois contact structure $\mathbb{S}^{5}\times\mathbb{T}^{2}$ with the $3$-dimensional $A_{k-1}$-singularity link as the binding, these non-contractible orbits in the binding become contractible in the Bourgeois contact manifold. Since $H_{1}(\Sigma(k,2,2);\mathbb{Z})$ is torsion, all Reeb orbits, including those non-contractible orbits, can be assigned with well-defined rational Conley-Zehnder indices [McL16, §4], which are at least $1$ by the computation in [MR18, §2.1]. In view of this, the 3-dimensional links of $A_{k-1}$-singularities are $-1$-ADC.
We next observe that the $k$-ADC condition is stable under the Bourgeois construction:
@@
If the binding of an open book $OB(\Sigma,\psi)$ is simply connected and $k$-ADC, then the Bourgeois contact manifold $BO(\Sigma,\psi)$ is $(k+2)$-ADC.
@@
In view of Lemma 2.4, we can assume that, for some fixed $\delta>0$ sufficiently small, on a neighborhood $B\times\mathbb{D}^{2}_{2\delta}\times\mathbb{T}^{2}$ of the (stabilized) binding of the open book, the Bourgeois contact form is
$\alpha_{BO}=h_{1}(r)\alpha_{B}+h_{2}(r)d\theta+\lambda_{std}$
where $\lambda_{std}$ is the standard Liouville form on the portion $\mathbb{D}^{2}_{2\delta}\times\mathbb{T}^{2}\simeq D^{*}_{2\delta}\mathbb{T}^{2}$ of unit cotangent bundle made of vectors of norm less than $2\delta$, and where the functions $h_{1},h_{2}$ are such that the following conditions hold:
  • $h_{1}(r)\equiv 1$ for $r\leq\delta$ and $h_{1}^{\prime}(r)<0$ otherwise;
  • $h_{2}(r)\equiv 0$ for $r\leq 3\delta/2$ and $h_{2}^{\prime}(r)>0$ otherwise.
Consider then the Hamiltonian $H(q_{1},p_{1},q_{2},p_{2})=\eta(r)(f(q_{1},q_{2})+g(r)),$ where $(q_{i},p_{i})$ are coordinates on $D_{2\delta}^{*}\mathbb{T}^{2}$, $r=\sqrt{p_{1}^{2}+p_{2}^{2}}$, $\eta$ is a bump cut-off function, $f\colon\mathbb{T}^{2}\to\mathbb{R}$ is a Morse-function on $\mathbb{T}^{2}$, and $g\colon[0,1]\rightarrow\mathbb{R}$ is Morse with a unique maximum at $r=0$. Following Bourgeois [Bou02a] we consider
$\alpha_{\epsilon}=(h_{1}+\varepsilon H)\alpha_{B}+\lambda_{std}$
on the $2\delta$-neighbourhood $B\times\mathbb{D}^{2}_{2\delta}\times\mathbb{T}^{2}$. Then one computes that the Reeb vector field is of the form
$\frac{1}{h_{1}+\varepsilon H-\lambda_{std}(X_{h_{1}}+\varepsilon X_{H})}\left[R_{B}-X_{h_{1}}-\varepsilon X_{H}\right],$
where we use the notation $X_{h}$ for the Hamiltonian vector field of a function $h:D^{*}\mathbb{T}^{2}\rightarrow\mathbb{R}$ with respect to $\omega_{std}=d\lambda_{std}$. Note that the flow generated by the vector field $X_{h_{1}}$ is just a rescaling of the geodesic flow on $D^{*}\mathbb{T}^{2}$ and is stationary precisely where $h_{1}^{\prime}=0$.
Now for a fixed action threshold $D$, taking $\varepsilon$ small, we see that the only contractible periodic orbits up to action $D$ must lie in the region where $h_{1}$ is constant, i.e. $B\times\mathbb{D}^{2}_{\delta}\times\mathbb{T}^{2}$. These orbits are of the form $\gamma_{q}=\gamma_{B}\times\{(q,r=0)\}$ with $q$ a critical point of $f$ and $\gamma_{B}$ a contractible Reeb orbit in the binding. Moreover, one can compute, using the properties of the Conley-Zehnder index (c.f. [Sal99, Section 2.4]) that
$\mu_{CZ}(\gamma_{q})=\mu_{CZ}(\gamma_{B})-2+\textrm{ind}_{q}(H)\geq\mu_{CZ}(\gamma_{B}),$
as $\mathrm{ind}_{q}(H)=\mathrm{ind}_{q}(f)+2\geq 2$. Thus the degree on $M^{2n-1}\times\mathbb{T}^{2}$ is
$|\gamma_{q}|=(n+1)-3+\mu_{CZ}(\gamma_{q})\geq(n-1)-3+\mu_{CZ}(\gamma_{B})+2=|\gamma_{B}|_{B}+2,$
where $|\gamma_{B}|_{B}$ denotes the degree of $\gamma_{B}$ as an orbit in $B$. Consider a decreasing sequence $\alpha_{B}^{i}$ on the binding whose contractible orbits have degrees at least $k$ as in the definition of $k$-ADC, i.e. up to some action threshold $D_{i}\to\infty$. By multiplying by appropriate constants we can assume that this sequence is strictly decreasing. We then multiply the adapted contact form on $OB(\Sigma,\psi)$ by an appropriate function $f_{i}<1$ so that the form on the binding $\alpha_{B}$ is replaced by $\alpha_{B}^{i}$ to obtain a decreasing sequence of forms $\alpha^{i}$ for the Bourgeois contact structure whose contractible orbits also have degrees at least $k+2$ up to the action threshold $D_{i}$. Then perturbing as above to obtain non-degeneracy near the binding, we obtain contact forms $\alpha_{\varepsilon_{i}}^{i}$ so that the action of the above contractible orbit $\gamma_{q}$ is $(1+\varepsilon_{i}H(q,0))$ times its $\alpha_{B}^{i}$-action and $\varepsilon_{i}\ll 1$ small.
Strictly speaking these contact forms determine distinct contact structures, as their kernels may differ, but applying Gray stability this can be remedied by a isotopy, which can be made arbitrarily close to the identity by shrinking $\varepsilon_{i}$ as needed. We deduce that the Bourgeois contact structure is $(k+2)$-ADC if the contact structure on $B$ is $k$-ADC. ∎
@@
When $B$ is not simply connected, there might be Reeb orbits of $B$ which are non-contractible in $B$ but contractible in the ambient open book $OB(\Sigma,\psi)$. These orbits should then be included in the verification of the ADC condition for the open book, even though they are not relevant for that of $B$. In our applications, this will only happen in settings where $H_{1}(B;\mathbb{Z})$ is torsion. In this case, if the contact structure on $B$ has torsion first Chern class, every Reeb orbit can be assigned with a well-defined rational Conley-Zehnder index [McL16, §4]. One can then consider the notion of $k$-ADC manifolds but for all Reeb orbits. The computation in Lemma 2.8 is still valid for the rational Conley-Zehnder index.
@@
Consider the Bourgeois contact manifold associated to the Milnor open book on $\mathbb{S}^{2n-1}$ coming from the $A_{k-1}$-singularity, whose binding is a Brieskorn sphere $B=\Sigma_{n-1}(k,2,\cdots,2)$ and whose page is the Milnor fiber $V_{n-1}(k,2,\cdots,2)=\{z_{0}^{k}+z_{1}^{2}+\dots+z_{n-1}^{2}=\epsilon\}\cap\mathbb{D}^{2n}$, for small $\epsilon>0$. When $n\geq 4$, $B$ is a $1$-ADC contact manifold according to Example 2.6 and simply connected. We thus obtain from Lemma 2.8 that the corresponding Bourgeois contact manifolds on $\mathbb{S}^{2n-1}\times\mathbb{T}^{2}$ are $1$-ADC. When $n=3$, the corresponding Bourgeois contact manifolds on $\mathbb{S}^{2n-1}\times\mathbb{T}^{2}$ are also $1$-ADC by Example 2.7 and Remark 2.9.
@@ 2.3. $\mathbb{S}^{1}-$invariant contact structures and convex decompositions
Bourgeois contact structures are $\mathbb{T}^{2}$-invariant, and therefore in particular $\mathbb{S}^{1}$-invariant. As discussed in [DG12, Section 6] any $\mathbb{S}^{1}-$invariant contact structure on a manifold $V^{2n}\times\mathbb{S}^{1}$ induces a decomposition of the first factor into so called ideal Liouville domains, as defined in [Gir20b]. In particular, we have a topological decomposition
@@
Any (constant) section $V\times\{pt\}$ is a convex hypersurface and the decomposition $V=V_{+}\cup\overline{V}_{-}$ is a convex decomposition. In the case of the Bourgeois contact manifolds the ideal Liouville pieces are given as products of the page of the initial open book with an annulus, i.e. $V_{\pm}=\Sigma\times\mathbb{D}^{*}\mathbb{S}^{1}$. For details related to the notion of ideal Liouville domains and to the Bourgeois case we refer to [DG12, MNW13, Gir20b]. The pieces in this decomposition are round contactizations of ideal Liouville domains and are referred to as Giroux domains following [MNW13].

### 14 Mar 2026 v5

@@ 2.2. ADC Bourgeois contact manifolds
In this section, we will assume that the first Chern class $c_{1}(\xi)$ of all contact structures is torsion. In this case, there is a well-defined (rational) Conley-Zehnder index $\mu_{CZ}(\gamma)$ for any non-degenerate contractible periodic Reeb orbit $\gamma$. If the contact manifold is of dimension 2n12n-1, then the SFT degree of $\gamma$ is defined as
$|\gamma|\coloneqq n-3+\mu_{CZ}(\gamma).$
Notice that this does not correspond to the degree in ($S^{1}$-equivariant) symplectic cohomology of $\gamma$ when viewed as a periodic Hamiltonian orbit, which is just given by $n-\mu_{CZ}$.
We now recall the following definition from [Zho21b], which generalizes the one from [Laz20a]. In what follows, if $\alpha_{1},\alpha_{2}$ are contact forms on a manifold $M$, we write $\alpha_{1}>\alpha_{2}$ if there exists a smooth function $f:M\rightarrow\mathbb{R}$ which satisfies $f>1$, such that $\alpha_{1}=f\alpha_{2}$.
@@
A contact structure $(M,\xi)$ is called $k$ asymptotically dynamically convex ($k$-ADC) if there is a sequence of contact forms $\alpha_{1}>\alpha_{2}>\cdots>\alpha_{i}>\dots$, and a sequence of real numbers $D_{i}$ with $D_{i}\to\infty$, so that all contractible periodic orbits of $\alpha_{i}$-action $D_{i}$ are non-degenerate and have degree [^1] From now on, unless specified otherwise, for simplicity degree will mean the SFT degree. >k>k.
In particular, 0-ADC as in [Zho21b] is just ADC as in [Laz20a]. An important special case is when there is a contact form $\alpha$ that is index-positive, meaning, following [CO18, Section 9.5], that all contractible periodic orbits are non-degenerate and have positive degree for $\alpha_{i}\equiv\alpha$ the given contact form. Here, we introduce a slight generalization of index-positivity as follows:
Definition 2.5.
A contact form $\alpha$ on a contact manifold $(Y,\xi)$ with $c_{1}(\xi)$-torsion, is called $k$-index positive, if the contact form $\alpha$ is Morse-Bott non-degenerate along contractible Reeb orbits and $|\gamma|:=n-3+\mu_{LCZ}(\gamma)$ is strictly bigger than $k$ for all contractible Reeb orbits $\gamma$, where $\mu_{LCZ}$ is the lower semicontinous extension of the Conley-Zehnder index.
We recall that the degree $|\cdot|$ in Definition 2.5 is known as the lower SFT degree, as introduced by McLean [McL16]. Recall also from [Bou02b, Definition 1.7] that the Reeb flow $\varphi_{t}$ of a contact form $\alpha$ is called Morse-Bott non-degenerate if closed periodic orbits with a given action $T$ come in closed smooth submanifolds $N_{T}$ so that the rank of ${\mathrm{d}}\alpha|_{N_{T}}$ is locally constant and $TN_{T}=\ker({\mathrm{d}}\varphi_{T}-\operatorname{Id})$ at all points of $N_{T}$.
We also point out that the index $\mu_{LCZ}$ is nothing else than the lowest Conley-Zehnder index of the Reeb orbits obtained from the Morse-Bott family after breaking the Morse-Bott symmetry using the perturbations in [Bou02b, Lemma 2.3]. Indeed, $\mu_{LCZ}(\gamma)$ equals $\mu_{RS}(\gamma)-\frac{1}{2}(\dim N_{T}-1)$ by [Bou02b, Lemma 2.4], where $\mu_{RS}$ is the Robbin-Salamon index [RS93] (it is called generalized Maslov index in [Bou02b, Section 5.2.2], see [Gut14] for a thorough discussion of all these indices); see also [McL16, §2]. By [Bou02b, Lemmas 2.3 and 2.4], $k$-index positive contact manifolds are $k$-ADC, as one can perturb the $k$-index positive Morse-Bott contact form into a shrinking sequence of contact forms satisfying the conditions in the definition of $k$-ADC.
A very useful aspect of the ADC condition is that it is preserved under flexible surgery [Laz20a], whereas index-positivity may not be. Lastly, we point out that the original definition of ADC in [Laz20a] used a non-increasing sequence of contact forms; however, one can always find a strictly decreasing sequence, as explained in [Laz20a, Remark 3.7.(4)].
@@
Consider the (2n1)(2n-1)-dimensional Brieskorn spheres $\Sigma_{n}(k,2,\cdots,2)$ given as the link of the $A_{k-1}$-singularity, i.e.
@@
This is an index-positive contact manifold. The degrees of the generators are at least 2n42n-4, provided that $n\geq 3$. This follows from [Ust99] in case $n=2m+1$ is odd, and from [vK08, Section 3] in general.
@@
for odd primes $p_{i}$, which according to [vK08, Section 4.1], is index positive in the Morse-Bott case as in Definition 2.5.
@@
For the case of $3$-dimensional Brieskorn manifolds, the computations are more involved as the manifolds are not simply connected. One can view links of $A_{k-1}$-singularities as quotients of the standard contact $3$-sphere so that contractible periodic orbits on the links can be lifted to the $3$-sphere. Those contractible orbits then have Conley-Zehnder indices at least $3$. The corresponding contact form on the link is Morse-Bott as it is the quotient of the round contact form on $S^{3}$, which is Morse-Bott. However, when we construct a $7$-dimensional Bourgeois contact structure $\mathbb{S}^{5}\times\mathbb{T}^{2}$ with the $3$-dimensional $A_{k-1}$-singularity link as the binding, these non-contractible orbits in the binding become contractible in the Bourgeois contact manifold. Since $H_{1}(\Sigma(k,2,2);\mathbb{Z})$ is torsion, all Reeb orbits, including those non-contractible orbits, can be assigned well-defined rational Conley-Zehnder indices [McL16, §4], which are at least $1$ by the computation in [MR23, §2.1]. In view of this, the 3-dimensional links of $A_{k-1}$-singularities are (1)(-1)-index positive.
We next observe that index-positivity is stable under the Bourgeois construction in the following sense:
@@
Assume that $\alpha$ is a contact form adapted to an open book decomposition with defining map $\Phi\colon M\to\mathbb{R}^{2}$ on $M$, and that its restriction $\alpha_{B}$ to the binding $B$ is $k$-index positive and $B$ is simply connected. Then, the Bourgeois form $\alpha_{BO}$ constructed as in Theorem 2.1 is $(k+2)$-index positive.
In particular, the associated Bourgeois contact structure on $M\times\mathbb{T}^{2}$ is $(k+2)$-ADC.
@@
An explicit computation using the expression in Lemma 2.3 shows that, on the neighborhood $\mathcal{N}=B\times\mathbb{D}^{2}\times\mathbb{T}^{2}$, the flow $\psi^{t}$ of $R_{BO}$ at time $t>0$ starting at a point $(b_{0},r_{0},\theta_{0},q_{1,0},q_{2,0})$ is given by
$b(t)=\psi_{B}^{t\mu(r_{0})}(b_{0})\,,\quad r(t)=r_{0}\,,\quad\theta(t)=\theta_{0}\,,\quad q_{1}(t)=q_{1,0}+t\nu(r_{0})\cos\theta_{0}\,,\quad q_{2}(t)=q_{2,0}-t\nu(r_{0})\sin\theta_{0}\,$,
where $\psi_{B}^{s}$ is the Reeb flow of $\alpha_{B}$ on $B$ at time $s$. Similarly, on the complement of $\mathcal{N}=B\times\mathbb{D}^{2}\times\mathbb{T}^{2}$, where $\nu=1$ and $\mu=0$, the flow at time $t$ is just given, in coordinates $(x,q_{1},q_{2})\in M\times\mathbb{T}^{2}$ and with starting point $(x_{0},q_{1,0},q_{2,0})$, by
$x(t)=x_{0}\,,\quad q_{1}(t)=q_{1,0}+t\cos\theta_{0}\,\quad q_{2}(t)=q_{2,0}-t\sin\theta_{0}\,.$
In particular, qualitatively, we have the following behavior. Closed Reeb orbits of $R_{BO}$ that touch $B\times\{0\}\times\mathbb{T}^{2}$ are in fact contained in it, and they organize in $\mathbb{T}^{2}$-parametric families, each coming from a closed Reeb orbit of $\alpha_{B}$. All the other closed Reeb orbits that lie in the complement of $B\times\{0\}\times\mathbb{T}^{2}$ project as a linear closed curve in the $\mathbb{T}^{2}$ factor. Hence, under our assumption of $B$ simply connected, the only contractible Reeb orbits of $R_{BO}$ are those of the form $\gamma_{q}=\gamma_{B}\times\{(0,q)\}$ with $\gamma_{B}$ a closed Reeb orbit of $\alpha_{B}$ and $q\in\mathbb{T}^{2}$.
Notice that the hypotheses on the function $h_{1}$ mean that $\nu(r)=\frac{2r}{h+r^{2}}$ for $r<1$, i.e. on $\mathcal{N}=B\times\mathbb{D}^{2}\times\mathbb{T}^{2}$, where $h:=h_{1}(0)>0$. Then if we use coordinates $(p_{1},q_{1},p_{2},q_{2})\in D^{*}\mathbb{T}^{2}\simeq\mathbb{D}^{2}\times\mathbb{T}^{2}\ni(p_{1},-p_{2},q_{1},q_{2})$ as in (1), we can rewrite the $q$-coordinates of the above described flow $\psi^{t}$ as
$q_{1}(t)=q_{1,0}+t\frac{p_{1,0}}{h+r_{0}^{2}},\quad q_{2}(t)=q_{2,0}+t\frac{p_{2,0}}{h+r_{0}^{2}}.$
It is then clear that the linearized flow along $B\times\{0\}\times\mathbb{T}^{2}$ splits in the $B$ direction and $D^{*}\mathbb{T}^{2}$ direction. The component in the $D^{*}\mathbb{T}^{2}$-direction of the linearized flow, that we denote by $\rho(t)$, is given, in the $(p_{1},q_{1},p_{2},q_{2})$-coordinates on $D^{*}\mathbb{T}^{2}$, by
$\rho(t)=\left[\begin{array}[]{cccc}1&0&0&0\\ \frac{t}{h}&1&0&0\\ 0&0&1&0\\ 0&0&\frac{t}{h}&1\end{array}\right].$
In particular, the linearized return map along $\gamma_{q}=\gamma_{B}\times\{(0,q)\}$ has, in the $D^{*}\mathbb{T}^{2}$-direction, as $1$-eigenspace the $\mathbb{T}^{2}$-direction (the zero section). Since $\alpha_{B}$ is Morse-Bott, it follows that $\alpha_{BO}$ is also Morse-Bott along $B\times\{0\}\times\mathbb{T}^{2}$. We also have
$\mu_{LCZ}(\gamma_{p})=\mu_{LCZ}(\gamma_{B})+\mu_{LCZ}(\{\rho(t)\mid 0\leq t\leq T\})$
where $T$ is the period of $\gamma_{p}$. Since the Robbin-Salamon index of the shear matrix path $\rho(t)$ is $1$ [Gut14, Lemma 26 (7)], we have $\mu_{LCZ}(\{\rho(t)\}_{0\leq t\leq T})=\mu_{RS}(\{\rho(t)\}_{0\leq t\leq T})-\frac{1}{2}\dim\ker(\rho(T)-\operatorname{Id})=0$. Thus the SFT degree on $M^{2n-1}\times\mathbb{T}^{2}$ is
$|\gamma_{q}|=(n+1)-3+\mu_{LCZ}(\gamma_{q})=(n-1)-3+\mu_{LCZ}(\gamma_{B})+2=|\gamma_{B}|_{B}+2\,.$
As $|\gamma_{B}|_{B}>k$ by assumption, we get $|\gamma_{q}|>k+2$.
Finally, the second conclusion is a direct consequence of the perturbation procedure in [Bou02b] as explained before the lemma. This concludes the proof. ∎
@@
When $B$ is not simply connected, there might be Reeb orbits of $B$ which are non-contractible in $B$ but contractible in the ambient open book $OB(\Sigma,\psi)$. These orbits should then be included in the verification of the ADC/index-positive condition for the open book, even though they are not relevant for that of $B$. In our applications, this will only happen in settings where $H_{1}(B;\mathbb{Z})$ is torsion. In this case, if the contact structure on $B$ has torsion first Chern class, every Reeb orbit can be assigned a well-defined rational Conley-Zehnder index [McL16, §4]. In this case, one can then consider the notion of $k$-ADC manifolds but for all Reeb orbits. The computation in Lemma 2.8 is still valid for the rational Conley-Zehnder index.
@@
Consider the Bourgeois contact manifold associated to the Milnor open book on $\mathbb{S}^{2n-1}$ coming from the $A_{k-1}$-singularity, whose binding is a Brieskorn sphere $B=\Sigma_{n-1}(k,2,\cdots,2)$ and whose page is Liouville homotopic to the Milnor fiber $V_{n-1}(k,2,\cdots,2)=\{z_{0}^{k}+z_{1}^{2}+\dots+z_{n-1}^{2}=\epsilon\}\cap\mathbb{D}^{2n}$, for small $\epsilon>0$. When $n\geq 4$, $B$ is a simply connected $1$-index positive contact manifold according to Example 2.6. We thus obtain from Lemma 2.8 that the corresponding Bourgeois contact manifolds on $\mathbb{S}^{2n-1}\times\mathbb{T}^{2}$ are $1$-ADC (and in fact even $3$-ADC, but $1$-ADC will be sufficient for our later purposes). When $n=3$, the corresponding Bourgeois contact manifolds on $\mathbb{S}^{5}\times\mathbb{T}^{2}$ are also $1$-ADC by Example 2.7 and Remark 2.9.
@@ 2.3. $\mathbb{S}^{1}$-invariant contact structures and convex decompositions
Bourgeois contact structures are $\mathbb{T}^{2}$-invariant, and therefore in particular $\mathbb{S}^{1}$-invariant. As discussed in [DG12, Section 6], any $\mathbb{S}^{1}$-invariant contact structure on a manifold $V^{2n}\times\mathbb{S}^{1}$ induces a decomposition of the first factor into so-called ideal Liouville domains, as defined in [Gir20b]. In particular, we have a topological decomposition
@@
Any (constant) section $V\times\{pt\}$ is a convex hypersurface, and the decomposition $V=V_{+}\cup\overline{V}_{-}$ is a convex decomposition. In the case of the Bourgeois contact manifolds, the ideal Liouville pieces are given as products of the page of the initial open book with an annulus, i.e. $V_{\pm}=\Sigma\times\mathbb{D}^{*}\mathbb{S}^{1}$. We refer to the existing literature for details about what just described: in [Gir20b] and [MNW13, §5.2] for ideal Liouville domains, in [DG12] for $\mathbb{S}^{1}$-invariant contact structures in general, and in [BGM22, §6.1, 6.3] for the case of Bourgeois structures as $\mathbb{S}^{1}$-invariant ones. The pieces in this decomposition are round contactizations of ideal Liouville domains and are referred to as Giroux domains following [MNW13].